OpenStudy (anonymous):

The Earth is approximately 150 million kilometers away from the Sun. If light travels at a speed of approximately 300,000 kilometers per second, about how long does it take for light from the Sun to reach the Earth?

OpenStudy (anonymous):

distance = rate times time time = distance divided by rate divide 150,000,000 by 300,000 to get your answer

OpenStudy (anonymous):

500

OpenStudy (anonymous):

\[\frac{1}{60}\frac{150000000}{300000}=\frac{25}{3}= 8+\frac{1}{3} \text{minutes}\]